OpenStudy (anonymous):

Prove the folliwing combinatorial identity if 1≤k

OpenStudy (anonymous):

\[1\le k <n\] \[\left(\begin{matrix}n-1 \\ k-1\end{matrix}\right) \left(\begin{matrix}n \\ k+1\end{matrix}\right) \left(\begin{matrix}n+1 \\ k\end{matrix}\right) = \left(\begin{matrix}n-1 \\ k\end{matrix}\right) \left(\begin{matrix}n \\ k-1\end{matrix}\right) \left(\begin{matrix}n+1 \\ k+1\end{matrix}\right)\]

OpenStudy (anonymous):

What do you want us to find? @zhengcl86

OpenStudy (anonymous):

this question says prove the folliwing combinatorial identity if 1≤k<n . this identity is known as the hexagon identity and relates terms in pascal's triangle

OpenStudy (anonymous):

then its got that whole equation on the bottom (n−1k−1)(nk+1)(n+1k)=(n−1k)(nk−1)(n+1k+1)

OpenStudy (anonymous):

This one is actually pretty easy if you just rewrite each combination as: \[\left(\begin{matrix}a \\ b\end{matrix}\right)=\frac{ a! }{ (a-b)!b! }\].